59,231.41J of heat is added to 1000 g of ice at -10 oC. How much ice melts?
Answer #1

given, M = 1000 g = 1 kg

we know,

specific heat of ice, c_ice = 2030 J/(kg K)

latent heat of fusion, Lf = 3.33*10^5 J/kg

let m is the of the ice that melts


now Apply,

Q = M*c*dT + m*Lf

59231.41 = 1*2030*(10-0) + m*3.33*10^5

==>m = (59231.41 - 20300)/(3.33*10^5)

= 0.117 kg or 117 grams <<<<<<<<<<<<<<<<<<<<<<<<------------Answer